OpenStudy (leahc246):

Verifying Identities: How would I make this expression equal tanx? sin^2x cosx ------ x ----- cosx 1

OpenStudy (leahc246):

I know tangent is sin/cos but I can't seem to get the algebra right

OpenStudy (acespeedfighter):

0.0000000000000000000000000000001% sure that it is not 1

OpenStudy (slenderman):

or you know you could cross multiply your duplicate and then divide what you have left

OpenStudy (leahc246):

That's what I was thinking... here's the original identity: tanx-sin(-x) ----------- 1 + cosx

OpenStudy (slenderman):

dear neptune

OpenStudy (leahc246):

Agreed

OpenStudy (slenderman):

how do you even.....

OpenStudy (triciaal):

sin 2x = 2 sinx cosx

OpenStudy (triciaal):

sin^2 x = 1 - cos^2 x

OpenStudy (triciaal):

what is the original problem?

OpenStudy (philomath):

\[\Large \frac{\tan(x)-\sin(-x)}{1+\cos(x)}\] \[\Large \frac{\tan(x)+\sin(x)}{1+\cos(x)}\] \[\Large \frac{\tan(x)+\sin(x)}{1+\cos(x)} * \frac{1-\cos(x)}{1-\cos(x)}\] \[\Large \frac{\tan(x)+\sin(x)-\tan(x)*\cos(x)-\sin(x)*\cos(x)}{1-\cos^2(x)}\] \[\Large \frac{\tan(x)+\sin(x)-\frac{\sin(x)}{\cos(x)}*\cos(x)-\sin(x)*\cos(x)}{\sin^2(x)}\] \[\Large \frac{\tan(x)-\sin(x)*\cos(x)}{\sin^2(x)}\] \[\Large \frac{\frac{\sin(x)}{\cos(x)}-\sin(x)*\cos(x)}{\sin^2(x)}\] \[\Large \frac{\sin(x)\left(\frac{1}{\cos(x)}-\cos(x)\right)}{\sin^2(x)}\] \[\Large \frac{\frac{1}{\cos(x)}-\cos(x)}{\sin(x)}\] \[\Large \frac{\frac{\cos^2(x)-1}{\cos(x)}}{\sin(x)} = \frac{\cos^2(x)-1}{\cos(x)} * \frac{1}{\sin(x)}\] \[\Large \frac{1-\sin^2(x)-1}{\cos(x)*\sin(x)} = \frac{\sin(x)*\sin(x)}{\cos(x)*\sin(x)} = \frac{\sin(x)}{\cos(x)} = \] tan(x)

OpenStudy (philomath):

Did you follow?

OpenStudy (phi):

if you change tan to sin/cos, the derivation is a bit shorter \[ \frac{\tan x + \sin x}{1+\cos x}= \frac{1}{1+\cos x}\left( \tan x +\sin x\right) \\ = \frac{1}{1+\cos x}\left( \frac{\sin x}{\cos x} +\sin x\right)\\ = \frac{1}{1+\cos x}\left( \frac{\sin x}{\cos x} + \frac{\cos x\sin x}{\cos x}\right)\\ = \frac{1}{1+\cos x}\left( \frac{\sin x(1+\cos x)}{\cos x} \right) \] cancel the 1+cos x to get sin /cos = tan